Find the slope of the normal to the curve X=1-asin theta, y= bcos^2 theta at theta=π/2.

Grade:12

1 Answers

Vikas TU
4 years ago
Dear student 
dx/dQ = -aCosQ 
y = bcos^2Q 
dy/dx = 2bCosQ(-SinQ)
= -2bSinQCosQ 
dy/dx = dy/dQ * dQ/dx  
= 2b/a SinQ 
Slope of normal at Q = pi/2 
= 1/(dy/dx) at Q = pi/2 
Slope o normal = -a/2b 
Good Luck